OpenStudy (anonymous):

x^2+y^2+12x+6y+36=0 find center and radius

hero (hero):

Re-write the equation of a circle in the form \((x - h)^2 + (y - k)^2 = r^2\)

OpenStudy (anonymous):

How so? \[x^2+y^2+12x+6y+36=0\]